Solve the system by the elimination method.
6x+7y=4
5x+8y=-1

Answer by mananth(16946)

6 x + 7 y = 4 .............1
Total value
5 x + 8 y = -1 .............2
Eliminate y
multiply (1)by -8
Multiply (2) by 7
-48 x -56 y = -32
35 x + 56 y = -7
Add the two equations
-13 x = -39
/ -13
x = 3
plug value of x in (1)
6 x + 7 y = 4
18 + 7 y = 4
7 y = 4 -18
7 y = -14
y = -2
x= 3
y= -2